If the automobile starter draws 150 A from the previous 12V battery (with stored energy of 1.73E6J once fully charged), how lplease explain, thanks!
Answer #1

relation between power, voltage and current is

P= VI = 12V(150A) = 1800W

P = E/t

t = E/P = 1.73* 10^6 J/1800 W = 961 s

option (B) is correct